# Grid Monitoring and Protection with Continuous Point-on-Wave Measurements and Generative AI

### v1

@@ 2.1 Wiener-Kallianpur Innovation Representation
By matching (almost surely) the autoencoder output $(\hat{X}_{t})$ with its input $(X_{t})$, Wiener-Kallianpur’s innovation representation is an ideal architecture for the compression and reliable communications of nonparametric random processes in the (Shannon) information-theoretic sense.
The idea of innovation representation became popular in engineering in the 1960s [15], especially in signal processing and control. The measurement-update step in Kalman filtering is an example of how innovations are used in estimation, prediction, and control. Despite the generality of the Wiener-Kallianpur innovation representation, its practical applications are limited to Gaussian and additive Gaussian models—the only two cases when causal autoencoder $(G,H)$ can be constructed explicitly and implemented in practice. No general computationally tractable way to extract the innovations sequence exists, even when the underlying probability distributions of $(X_{t})$ are known. Outside the recent work of [16], no existing work is capable of learning the innovation representation.
2.2 Learning Innovation Representation
In [16], the authors proposed a generative adversary network (GAN) based approach to learning the Wiener-Kallianpur innovation representation by jointly minimizing the Wasserstein distance between the latent process and the uniform IID process and the mean-squared error ($l_{2}$ distance) of the autoencoder output as an estimate of the input. The structure of the deep learning approach to innovation representation is shown in Fig. 3.
Figure 3: A Machine Learning Architecture for Innovation Representation.
The encoder $G_{\theta}$ and decoder $H_{\eta}$ are causal convolutional neural networks parameterized by coefficients $\theta$ and $\eta$, respectively. These parameters are optimized to match the autoencoder output $\hat{X}_{t}$ with its input $X_{t}$ (in the mean-squared sense) and impose constraints that the latent process $(V_{t})$ is an IID process. Here, we consider two possible types of innovations: the IID-uniform innovations as in [16] suitable for anomaly detection and IID-Gaussian for compression and communications. The latter is a slight but new variation designed to match the compression with the typical additive Gaussian characteristics of communication channels.
The learning algorithm of the autoencoder $(G_{\theta},H_{\eta})$ and discriminator $D_{\gamma}$ is of the GAN type with parameters $(\theta,\gamma,\eta)$ jointly optimized from the $l_{2}$-Wasserstein distance measures:
$L\big{(}(X_{t}),\theta,\eta\big{)}$ $:=\max_{\gamma}\Big{(}\mathbb{E}\Big{[}D_{\gamma}\big{(}(V_{t}),(U_{t})\big{)}\Big{]}+\lambda\mathbb{E}\Big{[}||(\hat{X}_{t})-(X_{t})||_{2}\Big{]}\Big{)}$ (2)
@@ 2.2 Learning Innovation Representation
In practice, finite-dimensional neural network input and finite-dimensional segments of random processes are used. In [16], structural convergence is established to show that the finite-dimensional implementation of the autoencoder converges to the theoretical infinite-dimensional innovation representation, as the dimensionality of the deep-learning architecture grows.
2.3 Anomaly and Novelty Detection
We present an innovation-based approach to anomaly and novelty detection, where the underlying probability models are unknown under normal and abnormal conditions. In particular, we consider a prototype hypothesis testing model involving random CPOW measurement ${\bm{X}}=(X_{t})$ either at a local IED, a substation, or at a control center:
${\cal H}_{0}:{\bm{X}}\sim P_{0}~{}~{}\mbox{vs.}~{}~{}{\cal H}_{1}:{\bm{X}}\sim Q\in\mathscr{Q}.$ (4)
The null hypothesis ${\cal H}_{0}$ represents data under normal operating conditions when CPOW measurement ${\bm{X}}$ is a random process following unknown probability distribution $P_{0}$. We assume that training data are available to learn an innovation representation of ${\bm{X}}$ discussed in Sec. 2 via offline or online learning. It is reasonable to assume that historical data under ${\cal H}_{0}$ are plentiful.
The alternative hypothesis ${\cal H}_{1}$ models the scenario when anomalous CPOW measurements follow an unknown distribution $Q\in\mathscr{Q}$. We do not specify the anomaly types, as they may not be enumerable. Nor do we assume that there are historical data to learn $\mathscr{Q}$ because, in practice, validated historical data for anomalies are rare.
The anomaly detection model (4) can be generalized for detecting new phenomena that are not necessarily faulty. Such a novelty detection problem can be modeled by
${\cal H}_{0}:{\bm{x}}\sim P\in\mathscr{P}~{}~{}\mbox{vs.}~{}~{}{\cal H}_{1}:{\bm{x}}\sim Q\in\mathscr{Q}.$ (5)
where $\mathscr{P}=\{P_{0},\cdots,P_{K}\}$ is a dictionary of observation classes identified and validated in the past, making hypothesis ${\cal H}_{0}$ composite. With autoencoders associated with each $P_{i}\in\mathscr{P}$ learned, the detection approach developed here can be applied.
Related Work
The detection model (4) belongs to the category of nonparametric hypothesis testing [17], for which classical approaches in statistics generally assume known probability model $P$ under ${\cal H}_{0}$. Standard techniques such as the Kolmogorov-Smirnov test and others [18] can be used for offline detection involving non-sequential data with a known model under ${\cal H}_{0}$.
Modern machine learning removes the assumption of known probability models under both ${\cal H}_{0}$ and ${\cal H}_{1}$, though techniques specializing in anomaly/novelty detection in time series are limited because the probability structures for nonparametric time series are infinite-dimensional. Standard approaches reduce the problem to a finite-dimensional approximation by considering finite-duration segments of time series samples. See, e.g., [19, 20, 21]. Such segmentations overlook the temporal dependencies of the time series, which are crucial in anomaly/novelty detection unless the segments are chosen to be very large. This results in a long detection delay and intractable computation complexity.
With an unknown probability model under ${\cal H}_{0}$, the state-of-the-art approaches roughly fall into three categories. One is based on classifiers such as the one-class support vector machine [22] and its many variants for different applications [23, 24, 25]. Such techniques implicitly assume that the support of the anomalous CPOW data distribution is largely disjoint from that of the anomaly-free. With only training data under ${\cal H}_{0}$, these techniques perform poorly when anomaly and anomaly-free models have overlapping support domains.
The second group of techniques simulates samples under ${\cal H}_{1}$ using hypothetical distributions some distance away from that under ${\cal H}_{0}$ [26, 27, 28]. It is, however, difficult to capture all possible alternative distributions. The third category captures many machine learning solutions developed in recent years, where the detection is performed in a transformed “feature domain”, where some types of confidence scores on the learned anomaly-free model are used to detect anomaly [29, 30, 31, 32]. As shown in [33], however, even with the perfect density estimate, detection may still perform poorly. Most relevant to our approach are the autoencoder-based machine-learning techniques for anomaly/novelty detection problems. See e.g., [34, 24, 25, 35, 36]. The use of independent component analysis in [35, 36] bears a structural resemblance to the innovation approach, with the main difference lying in the requirement of causality on the autoencoder and training in our approach. Graphical models have been proposed to exploit graphical structures exhibited in the training data under ${\cal H}_{0}$ [37, 38, 39]. Such techniques need to be better explored, especially in the context of detecting anomalies in time series.
With the phenomenal success of natural language processing, the idea of “transformer” has gained a large following, and there have been successful demonstrations recently in classification and some time series forecasting problems [40, 41]. The applicability of these techniques in power systems remains to be determined as power system data do not have the long-range dependencies typical natural language models, making the high-dimensional transformer-type of language models difficult to learn [42].
Detection via Innovation Representation
The approach presented here is based on the innovation representation discussed in Sec. 2 and was originally developed in [16, 43]. The most striking feature of the innovation-based anomaly detection is that the innovation autoencoder transforms the problem of anomaly detection with an unknown probability model under ${\cal H}_{0}$ to one under the known and simple model of IID uniform (or Gaussian). The new contribution here is the application of Neyman’s Smooth Test (NST) for anomaly detection and a sequential technique for quick detection.
Let $(V_{t})$ be the innovation sequence of the Wiener-Kaillianpur autoencoder $(G,H)$. The anomaly detection model (4) is transformed to the feature space of innovations
[…]
While the underlying hypothesis testing problem remains challenging because of the arbitrariness and the uncountable nature of the alternative probability models under ${\cal H}_{1}$, the innovation representation provides a bridge to the classical Goodness-of-Fit (GoF) problem, for which considerable literature exists [44].
Figure 4: Neyman’s Smooth Test (NST).
@@ 2.3 Anomaly and Novelty Detection
Neyman proposed the NST in 1937 [45] for testing (6) when the probability distribution under ${\cal H}_{1}$ is “smooth.” In the contemporary machine learning language, NST employs the orthogonal Legendre polynomial kernels $\{\pi_{k}\}$ as a kernel representation of probability distributions under ${\cal H}_{1}$. The schematic of NST is shown in Fig. 4 where inputs $(v_{t})$ are passed through Legendre polynomial kernels at the input layer. The output of the input layer is the sum of independent random variables at the Central-Limit Theorem scaling, resulting in a limiting Gaussian distributed $\sum_{i}y_{i}/\sqrt{N}$, and $\chi^{2}$ distributed $T({\bm{v}})$ with $K$-degree of freedom, asymptotically. The testing is a threshold decision based on
@@
NST was shown to be the locally most powerful test for uniformity testing [45], outperforming all other possible detection techniques in a neighborhood around the null hypothesis ${\cal H}_{0}$. In Rayner and Best [46], the author declared: “don’t use those [other] methods—use a smooth test!”
3 Compression and Data Analytic Architecture
This section presents an innovation-based compression and data analytic architecture that combines subband coding, innovation representation, and local/central anomaly/novelty detection. The schematic of the proposed compression system is shown in Fig. 5 with key components explained next.
Figure 5: A schematic of the innovation-based compression.
The upper branch of Fig. 5 is the schematic for an IED that takes analog current/voltage measurement $x(t)$. Power system signals are narrow passband signals with special subband spectral structure illustrated in Fig. 5. Around the fundamental frequency ±$\pm f_{0}$ is the fundamental frequency band regulated with frequency deviations within $W_{0}$. Typically $W_{0}<1$ Hz. Above the fundamental frequency band are harmonic subbands at the multiples of $f_{0}$, $f_{k}=kf_{0}$, that account for higher-order harmonics due to nonlinear devices in the network. There may also be inter-harmonic frequency components represented by the grey area. Higher-order harmonics and inter-harmonics may appear sporadically; they are modeled as on-off random processes.
Subband Coding
The A/D converter samples $x(t)$ at the Nyquist rate to obtain a digital CPOW signal $x[n]$. Depending on the spectral bandwidth of $x(t)$, the Nyquist sampling rate is typically in the tens of kHz (or higher) range, resulting in more than 600 K samples per cycle. The subband coding scheme proposed in [5] greatly reduces the Nyquist-sampled data rate via the subband decomposition and encodes each harmonic (and inter-harmonic) subband separately. By the passband Nyquist sampling theorem, a harmonic subband with passband bandwidth $W$ requires only a $2W$ sampling rate. For instance, if the fundamental subband has the bandwidth of $W_{0}=1$ (Hz), one sample per 30 cycles is sufficient to capture the signal in the fundamental frequency band fully. For an $m$ harmonic-subband spectrum, each with $W$ Hz passband bandwidth, $mW/30$ samples per cycle are sufficient for lossless compression. The subband coding block extracts the baseband components of $X(f)$, down-sampling the Nyquist-sampled data to achieve two to three orders of magnitude data rate reduction. The subband coding technique is well developed for multi-media compression and is embedded in practical audio/video communications data-streaming technologies. The subband encoder proposed in [5] can be easily implemented by off-the-shelf filter banks. See details in [5]. Note the per-cycle data samples of the subband CPOW signals $\{(x_{i}[n])\}$ are baseband representations of the signal spectrum, typically at the rate of less than one sample per cycle under normal operating conditions.
Innovation Autoencoder and Local Data Analytics
While the subband coding scheme significantly reduces the data rate of the direct Nyquist sampled $(x[n])$, further data rate reductions can be achieved through information-theoretic approaches that offer an optimal tradeoff between the compression rate and distortions of reconstruction. Additional data rate reduction can be achieved through local data analytics that adaptively selects harmonic subbands for transmission. For instance, harmonic and inter-harmonic subbands may be inactive under normal operating conditions. During significant network events, signal contents from only a few harmonic subbands may be informative.
The innovation autoencoder discussed in Sec. 2 is used to transform subband signals $\{(x_{i}[n])\}$ to a set of innovations sequences $\{(v_{i}[n])\}$, each is an IID uniform or Gaussian sequences. For minimum data compression over the typical additive Gaussian noise channel of the utility network, it is natural that the innovation sequence $(v_{i}[n])$ is IID Gaussian with matching mean and variance with $(x_{i}[n])$, from which optimal bits-allocation to each subband innovations can be determined. This is discussed as part of the vector quantization encoder in Sec. 3.
The local data analytics unit detects anomalies and novelties using techniques presented in Sec. 2.3. The local data analytics unit produces the local state sequence $(s_{t})$ that can be used to compress data further and provide summary statistics for the control center. Specifically, $(s_{t})$ may include the normal/fault state sequence and estimated fault locations when faults are detected locally. Under the normal system state, data from many subbands can be suppressed, and the control center can recall the suppressed signal if necessary.
Vector Quantization Encoder
Vector quantization converts innovations and state sequence $\{(v_{i}[n]),s[n])\}$ to bit-streams to be transmitted over the utility network (after multiplexing with bit streams from other IEDs.) The quantization schemes developed in [5] can be used. The new feature considered here is an information theoretic approach to compression not included in [5].
Consider the IID-Gaussian innovation sequences $(V_{i}[n])$, with mean and variance $(\mu_{i},\sigma_{i}^{2})$ for the $i$th subband innovations. Given the acceptable mean-square distortion of the reconstruction $D_{i}:=\mathbb{E}(\hat{X}_{i}[n]-X_{i}[n])^{2}$, the optimal compression compression is achieved by allocating distortion $D_{i}$ to subband $i$ constrained by $D=\sum_{i}D_{i}$ through the so-called inverse water-filling procedure [47]. The Gaussian rate-distortion function gives the minimum transmission rate subject to distortion measure
$R(D)=\sum_{i=1}^{m}\frac{1}{2}\log\frac{\sigma_{i}^{2}}{D_{i}}.$
Decoder at Substations or Control Center
The lower branch of Fig. 5 describes the building blocks of the decoding process at the control center. For the most part, this process is the inverse of the upper branch, following mostly the procedure described in [5]. The vector quantization decoder reproduces the innovation sequence $\{(\hat{v}_{i}[n])\}$ for each subband (for those subbands that are not suppressed.) The innovation autodecoder is used to produce the decompressed subband signal $\{(\hat{X}_{i}[n])\}$ (in the baseband form), and the subband decoder assembles the subband signal to form $\hat{x}[n]$ as the estimation of the original CPOW $x[n]$, for which the estimation error is controlled by $D$ in the rate-distortion function.
Central Data Analytics and Interactive Networking
Integrating data analytics with compression and networking provides the key component for interactive networking: the local IEDs can initiate the interaction based on local detection and transmit anomalous waveforms. Likewise, with the received data feeds from multiple IEDs, the central data analytics unit can initiate requests for high-resolution data when data feeds result in ambiguities and inference with low confidence. Here, we do not discuss networking protocols that realize such interactive networking.
4 Over-current Fault Detection and Protection
Traditional power system protection is based on a deterministic power system model. For short circuit faults, a pre-determined threshold (the pickup current) is set such that current flows exceeding the threshold trigger relay actions. The significant presence of distributed stochastic generations and IBRs causes two types of erroneous relay actions. One is the so-called protection blinding shown on the left panel of Fig. 6, where the stochastic distributed generation (SDG) offsets the high current flow through the protection device (PD), causing Type II (miss-detection) error in PD’s decision. The second type is the sympathetic tripping shown on the right panel of Fig. 6, where the PD experiences high current flow from SDG, causing a Type I (false alarm) error and a PD’s unnecessary tripping action.
Figure 6: Two types of erroneous protection actions caused by stochastic distributed generation. Left: Protection blinding. Right: Sympathetic tripping.
Related Work
Power system protection under stochastic and distributed generations is an active area of research because of the increasing integration of power electronics-based resources. In [49], a detailed characterization and classification of distribution system protection under bulk renewable energy sources is provided. Classical techniques are “hard-wired” that directly compare the measurements with fixed threshold [50]. Other methods include installing fault current limiter [51, 52] to mitigate the effect of SDG or view pickup current as the decision variables to an optimization program, taking into account network configurations [53, 54].
More sophisticated data-driven solutions have been proposed. Methods belonging to this category design pickup current settings corresponding to different DG output levels, and the DG conditions are usually classified via data-driven methods [55]. In [56], fuzzy logic is applied to determine the DG working states and transferred to relays to adjust their setting. The authors of [57] applied wavelet transformation to measurements and adopted a neural network to detect the Ferroresonance for the wind farm system. Relay settings can also be computed from estimated system parameters of a Thevenin equivalent system [58] or based on steady-state fault current [59]. Other data-driven methods are developed for networked sensors as a multi-agent system [60], assuming a network supporting communications among agents. The method proposed in [61] adjusts the over-current relay setting based on current measurements from relays downstream.
4.1 Innovation-Based Sequential Fault Detection
We propose an innovation-based sequential fault detection (ISFD) approach to reduce protection errors and decision time. ISFD is a data-driven approach that makes testing decisions based on the latent process of the innovation autoencoder trained with anomaly-free realizations of possibly multiple stochastic CPOW measurements. In contrast, conventional techniques typically compare current phasor magnitudes with a pre-determined threshold. Such approaches are vulnerable to stochastic distributed generations, resulting in high false positive and negative rates that lead to sympathetic tripping and protection blinding, respectively. ISFD uses empirical distribution-related statistics to separate normal from faulty measurements when, in the time domain, the waveform corresponding to the two cases is very close. This is an instance in which choosing the right feature for detection is crucial.
ISFD is designed to achieve short detection delay by implementing a sequential version of NST based on the classical notion of change detection. When distribution statistic shows that the distribution of observed CPOW has a large distance from the normal distribution $P_{0}$ under ${\cal H}_{0}$, the procedure terminates quickly by declaring fault. When the observed CPOW is close to $P_{0}$ under ${\cal H}_{0}$, the observed data will likely be from a normal waveform. No alert is necessary, but more data are collected to enhance detection accuracy.
ISFD implements the above sequential procedure using the so-called doubling search technique [62]. At each iteration, ISFD executes NST with a fixed-sized batch of samples. If the output of the detecting technique is ${\cal H}_{0}$, the batch size will be doubled, and the detecting technique will be re-run in the next iteration. An upper bound on the number of iterations is often set to avoid infinite detection delay. The detailed algorithm of the doubling search trick implementation of NST is presented in Algorithm 1. Once NST has made its detection decision, its decision is passed to the circuit breaker to carry out its protection action.
@@ 4.1 Innovation-Based Sequential Fault Detection
4.2 Performance Evaluations
Simulation setting
We evaluated the performance of ISFD and two competing methods using the IEEE 13-bus distribution network shown in Fig. 7. To account for randomness in the measurements introduced by renewable energy resources, we added a $1$MW DG controlled by real-world current profile collected at the micro-grid located in EPFL between node $692$ and $675$. The DG was driven by EPFL’s current profile at the battery connection [^1] Data can be found here: https://github.com/DESL-EPFL/Point-on-wave-Data-of-EPFL-campus-Distribution-Network, and current measurement samples were collected at $50$KHz sampling frequency. The entire system is simulated through MATLAB/Simulink.
Figure 7: IEEE 13 Bus System Configuration.
Fig. 7 shows multiple fault scenarios considered in the simulation study, with red arrows indicating fault locations and blue boxes for the protective relays. We simulated single-phase short circuit faults for each fault location, using MATLAB Simulink to generate CPOW measurements. Specifically, fault F1 and F3 were chosen to test the impact of SDG under sympathetic tripping whereas F2 was chosen to evaluate the system’s performance under protection blinding.
When a single-phase short circuit fault happens at F1, R3 is the primary relay for F1 and R2 a backup. During F1, sympathetic tripping at R4 may occur because the current feed from SDG downstream could increase significantly. Similarly, when F3 happens, R4 is also subject to sympathetic tripping due to current flow from downstream DG. In this case, the primary relay is R2 and the backup R1. On the other hand, the primary relay R5 for fault scenario F2 faces protection blinding because of DG’s support for current and voltage at R5 during fault at F2.
Algorithm Implementations and Performance Metrics

### v2

@@ 2.1 Innovation Representation Model
[…]
2.2 Learning IRM in Foundation Model Pre-training
IRM was introduced by Norbert Wiener, Gopinath Kallianpur, and Murray Rosenblatt in their seminal work [3, 4, 18, 24] and more recent work in high-dimensional statistics [25, 26]. Under the Wiener-Kallianpur (strong) innovation representation, there exists a causal autoencoder $({\cal G},{\cal H})$ that the encoder ${\cal G}$ maps causally the input $(x_{t})$ to an IID uniformly distributed innovation sequence $(\nu_{t})$, and the causal decoder ${\cal H}$ that maps $(\nu_{t})$ to $(\hat{x}_{t}\stackrel{{\scriptstyle\mbox{\tiny a.s.}}}{{=}}x_{t})$. Under the Rosenblatt (weak) innovation representation, the autoencoder input and output are matched in distribution $(\hat{x}_{t}\stackrel{{\scriptstyle\mbox{\tiny D}}}{{=}}x_{t})$.
Fig. 3 shows the schematic pre-training of the IRM, first proposed in [27]. The innovation autoencoder is defined by a causal convolution neural network (CNN) encoder ${\cal G}_{\theta}$ with parameter $\theta$ and causal CNN decoder ${\cal H}_{\eta}$ with parameter $\eta$. These parameters are optimized to match the autoencoder output $\hat{x}_{t}$ with its input $x_{t}$ (in the mean-squared sense) and impose the constraint that the latent process $(v_{t})$ is IID.
Figure 3: A Machine Learning Architecture for Innovation Representation Model.
Specifically, the objective function that generates stochastic gradient descents combines the Wasserstein distance error produced by the innovation discriminator $D_{\gamma}$ and the mean-squared decoding error between the input $(x_{t})$ and output $\hat{x}_{t}$ of the autoencoder. The joint optimization of the neural network parameters $(\theta,\eta,\gamma)$ is given by
$L\big{(}\theta,\eta\big{)}$ $:=\max_{\gamma}\Big{\{}\tilde{\mathbb{E}}\Big{[}D_{\gamma}\big{(}(v^{(\theta)}_{t}),(u_{t})\big{)}\Big{]}+\lambda\tilde{\mathbb{E}}\Big{[}||(\hat{x}^{(\theta,\eta)}_{t})-(X_{t})||_{2}\Big{]}\Big{\}}$ (1)
@@ 2.2 Learning IRM in Foundation Model Pre-training
where $\tilde{\mathbb{E}}$ is the sample average operator. Note that the $l_{2}$ norm penalty on the decoding error used in (1) aims to achieve a (strong) mean-squared error reconstruction of the input by the decoder in the Wiener-Kallanpur representation. For the Rosenblatt relaxation of the innovation representation is used, the $l_{2}$ norm $\mathbb{E}\big{[}||(\hat{X}_{t})-(X_{t})||_{2}\big{]}$ is replaced by the Wasserstein distance measure $W\big{(}(\hat{X}_{t}),(X_{t})\big{)}$.
The above optimization can be realized using standard machine learning library [28, 29]. In practice, pre-training operates on finite-dimensional inputs and segments of random processes, with finite neural network architectures. In [27], the authors established structural convergence, demonstrating that the finite-dimensional implementation converges to the theoretical infinite-dimensional innovation representation as the architecture’s complexity increases.
3 Adaptation for Power System Anomaly Detection
We present the foundation model adaptation for anomaly detection by formulating a prototype hypothesis testing model involving stochastic WMU measurements ${\bm{x}}=(x_{t})$:
${\cal H}_{0}:{\bm{x}}\sim P_{0}~{}~{}\mbox{vs.}~{}~{}{\cal H}_{1}:{\bm{x}}\sim Q\in\mathscr{Q},$ (3)
where $P_{0}$ is the probability model for WMU time series ${\bm{x}}$ under normal operating conditions, and $\mathscr{Q}$ includes all abnormal probability distributions.
The above model (3) can be extended for novelty detection, where the objective is to catalog encountered abnormalities in an anomaly dictionary and detect whether the WMU measurements represent a new type of anomaly not encountered before. This generalization can be formulated as:
${\cal H}_{0}:{\bm{x}}\sim P\in\mathscr{P}~{}~{}\mbox{vs.}~{}~{}{\cal H}_{1}:{\bm{x}}\sim Q\in\mathscr{Q},$ (4)
where $\mathscr{P}={P_{0},\dots,P_{K}}$ is the dictionary of WMU measurement classes previously identified and validated either as known anomalies or as data under normal operating conditions.
In practice, neither $P_{0}$ nor $\mathscr{Q}$ is known, and the hypothesis testing formulations (3,4) are ill-defined. It is the IRM that transforms the ill-defined problem to a well-defined classic hypothesis testing problem. We assume that large data sets exist for measurements under normal operating conditions, enabling the pre-training of the IRM of the foundation model. Such an assumption is reasonable in practice and necessary for all AI foundation models. With a pre-trained autoencoder, the IRM-based foundation model transforms (3) to a challenging but tractable Goodness-of-Fit test problem:
${\cal H}^{\prime}_{0}:\hbox{\boldmath$\nu$\unboldmath}\sim P^{\prime}_{0}~{}~{}\mbox{vs.}~{}~{}{\cal H}^{\prime}_{1}:\hbox{\boldmath$\nu$\unboldmath}\sim Q\in\mathscr{Q}^{\prime}.$ (5)
where $P^{\prime}_{0}$ corresponds to the IID-uniformly distributed innovation sequence (under normal operating conditions), and $\mathscr{Q}^{\prime}$ is the collection of probability models for which the latent process of the foundation model encoder is not IID-uniform.
3.1 Related Work
The hypothesis testing model in (3) falls within the category of nonparametric composite hypothesis testing [31], where classical statistical approaches typically assume a known probability model $P$ under ${\cal H}_{0}$. Standard techniques, such as the Kolmogorov-Smirnov test and related methods [31], are well-suited for offline detection involving non-sequential data under ${\cal H}_{0}$. However, these approaches cannot apply to cases when probability models are unknown or when sequential data are involved.
Modern machine learning approaches alleviate the assumption of known probability models under both ${\cal H}_{0}$ and ${\cal H}_{1}$. Despite this, anomaly and novelty detection in time series remains challenging due to the infinite-dimensional nature of nonparametric time series probability structures. Standard methods address this by approximating the problem in a finite-dimensional space, often segmenting the time series into finite-duration samples (see, e.g., [32, 33, 34]). However, such segmentations often neglect temporal dependencies, which are critical for accurate detection. Large segments can mitigate this issue but result in long detection delays and prohibitive computation costs.
When the probability model under ${\cal H}_{0}$ is unknown, state-of-the-art detection techniques can be grouped into three broad categories:
  1. 1.
    Classifier-based approaches: Methods such as one-class support vector machines (SVM) [35] and their variants [36, 37, 38] implicitly assume that the support of the anomalous data distribution is largely disjoint from the anomaly-free distribution. While these techniques are effective in some cases, they perform poorly when the anomaly and anomaly-free models have overlapping support domains.
  2. 2.
    Simulations with hypothetical anomalies: These methods simulate samples under ${\cal H}_{1}$ using hypothetical distributions that are some distance from $P_{0}$ [39, 40, 41]. However, capturing all possible alternative distributions is inherently challenging, resulting in performance gaps caused by anomalies not covered by simulations.
  3. 3.
    Feature-domain confidence scores: Many recent machine learning solutions operate in a transformed feature domain, using confidence scores on a learned anomaly-free model to detect anomalies [42, 43, 44, 45]. As shown in [46], even with a perfect density estimate, these methods may struggle to achieve high detection performance. Autoencoder-based techniques are particularly relevant here [47, 37, 38, 48, 49].
The advent of LLM with a transformer architecture and attention mechanisms has shown promise in classification and time series forecasting tasks [50, 51]. However, the applicability of transformer-based architectures for anomaly detection in power systems is not well understood. Power system data often lack the long-range dependencies characteristic of natural language, making transformer models challenging to train effectively [23].
3.2 Foundation model Adaptation for Anomaly Detection
@@
The Neyman’s Smooth Test (NST) was first proposed in 1937 [53] for testing IID-uniform samples against the alternative distribution for non-IID uniform samples as in (5). In the contemporary machine learning language, NST can be viewed as a kernel-based classifier using the orthogonal Legendre polynomial $\{\pi_{k}\}$ as a kernel representation of probability distributions under ${\cal H}_{1}$.
@@
The schematic of NST is shown in Fig. 4 where the innovation input $(v_{t})$ are passed through Legendre polynomial kernels at the input layer. The output of the input layer is the sum of independent random variables at the Central-Limit Theorem scaling, resulting in a limiting Gaussian distributed $\sum_{i}y_{i}/\sqrt{N}$, and $\chi^{2}$ distributed $T({\bm{v}})$ with $K$-degree of freedom, asymptotically. The testing is a threshold decision based on
@@
NST was shown to be the locally most powerful test for uniformity [53], outperforming all other possible detection techniques in a neighborhood around the null hypothesis ${\cal H}_{0}$. In [54], Rayner and Best declared: “don’t use those [other] methods—use a smooth test!”
NST was originally designed for fixed-sample size testing. When testing time series where measurements arrive sequentially, a sequential test can result in a considerable speed-up in detection time. Here, we propose a sequential NST (SNST) to reduce detection errors and decision time. SNST is a data-driven approach that makes testing decisions using the latent innovation sequence of the innovation encoder. It is designed to achieve short detection delay by implementing a sequential version of NST, following the framework of change detection. When distribution statistic shows that the distribution of $(\nu_{t})$ exhibits significant deviation from the IID-uniform characteristics, the procedure terminates quickly by declaring fault. When $(\nu_{t})$ remains close to IID-uniform, no alert is made as more data samples are collected to enhance detection accuracy.
@@
SNST implements the above sequential procedure using the so-called doubling search technique [55]. At each iteration, SNST executes NST with a fixed-sized batch of samples. If the output of the detecting technique is ${\cal H}_{0}$, the batch size will be doubled, and the detecting technique will be re-run in the next iteration. An upper bound on the number of iterations is often set to avoid infinite detection delay. The detailed algorithm of the doubling search trick implementation of NST is presented in Algorithm 1. Once NST has made its detection decision, its decision is passed to the circuit breaker to carry out its protection action.
4 Adaptation for Over-Current Protection
Traditional power system protections are based on a deterministic power system model. For short-circuit faults, the standard practice is to pre-set the threshold (the pickup current) such that current flows exceeding the threshold trigger protection relay actions.
The significant presence of stochastic distributed generation (SDG) and IBRs causes two types of erroneous relay actions shown in Fig. 5. One is the so-called protection blinding shown on the left panel, where SDG offsets the high current flow through the protection device (PD), causing a Type II (miss-detection) error in the protection decision. The other type is the sympathetic tripping shown on the right panel, where the high current flow from the SDG causes the Type I (false alarm) detection error that triggers protection action when it should not be activated.
Figure 5: Two types of erroneous protection actions caused by stochastic distributed generation (SDG). Left: Protection blinding. Right: Sympathetic tripping.
4.1 Related Work
Power system protection under significant SDG is an active area of research because of the increasing integration of power electronics-based resources. In [56], a detailed characterization and classification of distribution system protection under bulk renewable energy sources is provided. Classical techniques typically compare the measurements with pre-fixed thresholds [57]. Other methods include installing fault current limiter [58, 59] to mitigate the effect of SDG or treating pickup current as a decision variable in an optimization program, taking into account network configurations [60, 61].
More sophisticated data-driven solutions have been proposed. Methods belonging to this category set current threshold settings based on either estimated SDG levels [62, 63, 64]. Relay settings can also be computed from estimated system parameters of a Thevenin equivalent system [65], steady-state fault current [66], and state estimation [67]. Other data-driven methods are developed for networked sensors as a multi-agent system [68], assuming a network supporting communications among agents. The method proposed in [69] adjusts the over-current relay setting based on current measurements from relays downstream.
4.2 Adaptation for Data-driven Over-current Protection
We propose an IRM-based foundation model approach, adapting to the sequential anomaly detection approach described in Sec. 3 for data-driven over-current protection in the presence of significant SDG, where WMU measurements are used for detecting over-current faults.
Figure 6: Schematic of the foundation model adaptation for over-current protection. Left panel: the IEEE 13 bus distribution system where protection relays R1 to R5 are labeled by the squares. Right panel: the implementation of SNST for over-current fault detection.
Fig. 6 shows a schematic for the foundation model adaptation for over-current protection. The left panel shows the IEEE 13 bus distribution system where protection relays R1-R5 are deployed. Short-circuit faults may occur at various locations, as illustrated by F1 to F3. With SDG injecting stochastic renewable power between bus 692 and 675, it will impact the primary relay action of R5, secondary relay R4, and possibly others.
[…]
4.3 Baseline Comparisons